OpenStudy (anonymous):

Simplify: (9 – 6)2 × (6 – 6 3)

OpenStudy (anonymous):

hmm, question. is it 63 at the end of the second bracket? or is 3 outside the brackets?

OpenStudy (anonymous):

If your question is "(9-6)(2) x (6-63)"... First deal with what is inside the brackets: (9-6)(2) x (6-63) = (4)(2) x (-57) and then multiply. (4)(2) x (-57) = (8) x (-57) = -456 So your answer would be -456. If your question is "(9-6)(2) x (6-6)(3)" Then again, we first deal with what is inside the brackets: (9-6)(2) x (6-6)(3) = (4)(2) x (0)(3) We immediately see that we will be multiplying 0 by everything so our answer is just 0.